OpenStudy (anonymous):

4x+17>and 3x-19>-13 solve

OpenStudy (anonymous):

somthing is wrong in question

OpenStudy (anonymous):

4x+17>5

OpenStudy (anonymous):

\[\ge\]

OpenStudy (anonymous):

this symbol goes instead of the one i entered

OpenStudy (anonymous):

\[4x+17\ge 5 and 3x-19\ge -13 solve\]

OpenStudy (anonymous):

\[\text{Reduce}[4x+17\geq 5\&\& 3x-19\geq -13] \text{ is}\text{ }x\geq 2 \]